OpenStudy (anonymous):

What is the value of the expression 8^–2? A. -16 B. 1/64 C. 1/16 D. 64

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

When something is raised to a negative exponent, you can write it as a reciprocal. \[a^{-b} = \frac{1}{a^{b}}\]

OpenStudy (anonymous):

so the answers could be B or C

OpenStudy (ipwnbunnies):

Correct. Do you know what 8^2 is?

OpenStudy (ipwnbunnies):

\[8^{-2} = \frac{1}{8^{2}}\]

OpenStudy (anonymous):

1/8^2 = 1/64

OpenStudy (anonymous):

Thanks! everyone!